Allan is explaining the wonders of nature to his son. He and his son went to a roof top of a 120m-high building, then he threw a stone from the rooftop.



1. Neglecting the air friction, what is the acceleration of the stone due to gravity?

2. At what time will the stone fall to the ground?

3. With what velocity will a stone strike the ground if dropped from the top of a

building?





1
Expert's answer
2021-06-21T16:33:25-0400

1. acceleration of the stone due to gravity "\approx" 9.81 m/s2

2. "h={1\over 2}gt^2"


"\implies 120={1\over 2}\times 9.81\times t^2"

"\implies t=\sqrt{120×2\over 9.81}=4.9462s"

3. V=gt =9.81×4.9462=48.5222 m/s